Calculate the pH and percent dissociation of 0.25 M NH3, Kb = 1.8 × 10^–5
Answer #1

Hucu +x +x NH3 (a) - Nha) - OH() 0.25 m -x [ 0.25-8 Kb = [ NHI [ CHO [ NH3] 1-8810-5 = x.x 0.25-8 2= 0.00212 M Pescend dissoc